Gabriela is 2 times as old as Brandon. Thirty-five years ago, Gabriela was 7 times as old as Brandon.
How old is Brandon now?

Answer by Maths68(1474)

Let
Present age of Gabriela = g
Present age of Brandon = b


Gabriela is 2 times as old as Brandon
g=2b..........(1)


Thirty-five years ago
Gabriela was = g-35
and
Brandon was = b-35

Then

Gabriela was 7 times as old as Brandon
g-35=7(b-35)
g-35=7b-245
g=7b-245+35
g=7b-210.............(2)

Put the value of g from (1) to (2)
2b=7b=-210
2b-7b=-210
-5b=-210
Divide both sides of above equation by -5
-5b/-5=-210/-5
b=42
Put the value of b in (1)
g=2b
g=2*42
g=84


Present age of Gabriela = g = 84 years old
Present age of Brandon = b = 42 years old